40m QIL Cryo_Lab CTN SUS_Lab TCS_Lab OMC_Lab CRIME_Lab FEA ENG_Labs OptContFac Mariner WBEEShop
  40m Log  Not logged in ELOG logo
Entry  Fri Oct 16 10:15:02 2009, steve, Update, IOO, IP ang & pos recentered ip4d.jpg
    Reply  Thu Oct 22 08:45:58 2009, steve, Update, IOO, IP ang & pos recentered pointing1000.jpg
Message ID: 2102     Entry time: Fri Oct 16 10:15:02 2009     Reply to this: 2132
Author: steve 
Type: Update 
Category: IOO 
Subject: IP ang & pos recentered 

Pointing stability of 4 days. Initial pointing does not go through suspended optics. It is launched  right after the Piezo Jena steering mirrors in the BS chamber.

IP-ANG on epics screen is  C1:ASC-IBQPD_X and Y in dataviewer  were recentered. This beam is clipping a bit in ETMX chamber  pick off mirror.

IP-POS pick  off is in the BS chamber and it's qpd on the BS_ISCT This beam is also clipping just a little bit. This is easy to fix. We'll have to remove an iris from the BS optical levers table.

note: arms were not locked when I recentered

Attachment 1: ip4d.jpg  93 kB  | Hide | Hide all
ip4d.jpg
ELOG V3.1.3-